# Structure-Enhanced Protein Instruction Tuning: Towards General-Purpose Protein Understanding

### v2

@@ B.4 More ablation studies
B.5 More case studies
More visualization for workload of experts in SEPIT.
@@ B.5 More case studies
Figure 5: Workload of experts in SEPIT for protein tokens and text tokens.
Figure 6: Workload of each expert in SEPIT for protein tokens and text tokens.
More cases for general-purpose protein understanding ability of SEPIT.
More cases of comparison is shown in Table 14.
Table 14: More cases on general-purpose protein understanding ability of SEPIT.
[Uncaptioned image] Q7CS13 (Swiss-Prot) Question: What is the protein family of <protein>? Ground Truth: Belongs to the metallo-dependent hydrolases superfamily. Atu3266/EF_0837 deacetylase family. Response from SEPIT-TinyLlama-MoEs: Belongs to the metallo-dependent hydrolases superfamily. Response from PIT-TinyLlama-MoEs: Belongs to the metallo-dependent hydrolases superfamily. Type II family. Response from Llama-Chat: Belongs to the metallo-dependent hydrolases superfamily. HutI family. Response from GPT-4-turbo: To understand the biological processes of the provided protein sequence, one must first consider its origin, functions, and structure. However, without direct access to databases or running bioinformatic tools right now… [Uncaptioned image] 6me7 (RSDB PDB) Question: Does this protein have cis-peptide linkages,<protein>? Ground Truth: No Response from SEPIT-TinyLlama-MoEs: No Response from PIT-TinyLlama-MoEs: Yes Response from Llama-Chat: Yes Response from GPT-4-turbo: No
More cases of our proposed SEPIT-TinyLlama-MoEs:
@@
    • Instruction: What are the short sequence motifs that are present in <protein>?
@@
    • Instruction: What is the function of <protein>?
@@
    • Instruction: Give me the function caption of <protein>.
@@
    • Instruction: Which protein family does <protein> belong to?
@@
    • Instruction: Which protein family does <protein> belong to?
@@
    • Instruction: What is the closest related protein family for <protein>?
@@
    • Instruction: What molecular function is associated with <protein>?
@@
    • Instruction: What cellular component is associated with <protein>?

### v3

@@ C.3. More Ablation Studies
C.4. More Case Studies
More Visualization for Workload of Experts in SEPIT
@@ C.4. More Case Studies
Figure 5. Workload of experts in SEPIT for protein tokens and text tokens.
Figure 6. Workload of each expert in SEPIT for protein tokens and text tokens.
More Cases for General-Purpose Protein Understanding Ability of SEPIT
More examples of our proposed SEPIT-TinyLlama-MoEs are shown below:
@@
    • Instruction: What are the short sequence motifs that are present in ¡protein¿?
@@
    • Instruction: What is the function of ¡protein¿?
@@
    • Instruction: Give me the function caption of ¡protein¿.
@@
    • Instruction: Which protein family does ¡protein¿ belong to?
@@
    • Instruction: Which protein family does ¡protein¿ belong to?
@@
    • Instruction: What is the closest related protein family for ¡protein¿?
@@
    • Instruction: What molecular function is associated with ¡protein¿?
@@
    • Instruction: What cellular component is associated with ¡protein¿?
@@
C.5. More Analysis
Analysis of Error Patterns
First, we present cases of SEPIT-TinyLlama-MoEs under major problem categories, with underlined text indicating differences from the ground truth (note that these differences are not necessarily errors). For each type, we attempt to demonstrate possible error patterns that might occur.
Function
  • Instruction: Explain the function of ¡protein¿. (A5FQQ5)
    • Response: This protein promotes the GTP-dependent binding of aminoacyl-tRNA to the A-site of ribosomes during protein biosynthesis.
    • Ground Truth: This protein promotes the GTP-dependent binding of aminoacyl-tRNA to the A-site of ribosomes during protein biosynthesis.
  • Instruction: Give me the function caption of ¡protein¿. (P44603)
    • Response: Catalyzes the transfer of an acyl group from acyl-phosphate (acyl-PO(4)) to glycerol-3-phosphate (G3P) to form lysophosphatidic acid (LPA). This enzyme utilizes acyl-phosphate as fatty acyl donor, but not acyl-CoA or acyl-ACP.
    • Ground Truth: Catalyzes the transfer of an acyl group from acyl-phosphate (acyl-PO(4)) to glycerol-3-phosphate (G3P) to form lysophosphatidic acid (LPA). This enzyme utilizes acyl-phosphate as fatty acyl donor, but not acyl-CoA or acyl-ACP.
  • Instruction: Can you define the function profile of the ¡protein¿? (B0BN86)
    • Response: Plays an essential role in mitochondrial morphogenesis.
    • Ground Truth: Plays a role in mitochondrial morphogenesis.
Similarity
  • Instruction: Which protein family does ¡protein¿ belong to? (Q99PZ1)
    • Response: Belongs to the 4-hydroxy-2-oxovalerate aldolase family.
    • Ground Truth: Belongs to the 4-hydroxy-2-oxovalerate aldolase family.
  • Instruction: Which protein class does ¡protein¿ fall into? (A8IQE0)
    • Response: Belongs to the CCDC40 family.
    • Ground Truth: Belongs to the CCDC39 family.
  • Instruction: To which protein family ¡protein¿ is classified? (P58600)
    • Response: Belongs to the UDP-N-acetylglucosamine-2-epimerase family.
    • Ground Truth: Belongs to the UDP-N-acetylglucosamine 2-epimerase family.
Subcellular location
  • Instruction: Where is ¡protein¿ located in the cell? (B0KZ78)
    • Response: Loosely associated with the inner membrane.
    • Ground Truth: Loosely associated with the inner membrane.
  • Instruction: Could you describe the subcellular location of ¡protein¿? (Q5M7N9)
    • Response: Localizes to endoplasmic reticulum-plasma membrane contact sites (EPCS). Recruited to the cell surface via the third C2 domain.
    • Ground Truth: Localizes to endoplasmic reticulum-plasma membrane contact sites (EPCS). Recruited to the cell membrane via the third C2 domain.
  • Instruction: What is the subcellular location of ¡protein¿? (Q9BXH1)
    • Response: Localized to the mitochondria in order to induce cytochrome c release.
    • Ground Truth: Localized to the mitochondria in order to induce cytochrome c release.
Induction
  • Instruction: Describe the effects of environmental factors of ¡protein¿’s expression. (P66808)
    • Response: By heat shock.
    • Ground Truth: By heat shock at 50 degrees Celsius.
  • Instruction: What are the environmental factors that induce the expression of ¡protein¿? (P83404)
    • Response: By bacterial infection.
    • Ground Truth: By bacterial and parasitic hemolymph and gut infections.
  • Instruction: What environmental factors causes the upregulation of ¡protein¿? (D2TDA7)
    • Response: By DNA damage, as part of the SOS response.
    • Ground Truth: By DNA damage, as part of the SOS response.
Molecular Function/Biological Process/Cellular Component (Gene Ontology)
  • Instruction: Which GO molecular function terms have ¡protein¿ been assigned to? (Q7VY99)
    • Response: transaldolase activity
    • Ground Truth: transaldolase activity
  • Instruction: What biological process is associated with ¡protein¿? (Q6PDU1)
    • Response: mRNA processing; regulation of alternative mRNA splicing, via spliceosome; RNA splicing
    • Ground Truth: mRNA splicing; regulation of alternative mRNA splicing, via spliceosome; response to vitamin E
  • Instruction: Which GO terms outline the cellular components of ¡protein¿? (F4HZB9)
    • Response: plasma membrane; side of membrane
    • Ground Truth: plasma membrane; side of membrane
Short Sequence Motif
  • Instruction: Can you identify and list all the motifs that are predicted to be present in ¡protein¿? (Q83I16)
    • Response: ’HIGH’ region; ’KMSKS’ region
    • Ground Truth: ’HIGH’ region; ’KMSKS’ region
  • Instruction: What are the short sequence motifs that are predicted to be present in ¡protein¿? (Q5R4H0)
    • Response: LXXLL motif; Methionine segments (Mets) motif
    • Ground Truth: LXXLL motif
  • Instruction: What are the short sequence motifs that are present in ¡protein¿? (B7H6M7)
    • Response: PRPP-binding
    • Ground Truth: PRPP-binding
Tissue Specificity
  • Instruction: In which tissues is the expression of ¡protein¿ absent? (Q5XIJ5)
    • Response: Expressed in brain (at protein level).
    • Ground Truth: Expressed in brain (at protein level). Expressed in hippocampal neurons.
  • Instruction: Describe the tissue-specific expression pattern of ¡protein¿? (Q6R2K3)
    • Response: Expressed in roots, leaves, stems, flowers and siliques.
    • Ground Truth: Expressed in seedlings, roots, stems, leaves, flowers and siliques.
  • Instruction: What is the tissue-specific expression pattern of ¡protein¿? (C0HJA8)
    • Response: Expressed by the venom duct.
    • Ground Truth: Expressed by venom duct.
Activity Regulation
  • Instruction: Describe the activity regulatory mechanism of ¡protein¿ associated enzymes, transporters, microbial transcription factors. (A3MKU1)
    • Response: Activated by UTP.
    • Ground Truth: Inhibited by UTP.
  • Instruction: What is the activity regulatory mechanism of ¡protein¿? (Q9US43)
    • Response: Allosterically activated by GTP.
    • Ground Truth: Allosterically activated by GTP.
  • Instruction: Tell me about the activity regulatory mechanism of ¡protein¿. (Q63486)
    • Response: The activation of GTP-binding proteins is generally mediated by a guanine exchange factor (GEF), while inactivation through hydrolysis of bound GTP is catalyzed by a GTPase activating protein (GAP).
    • Ground Truth: The activation of GTP-binding proteins is generally mediated by a guanine exchange factor (GEF), while inactivation through hydrolysis of bound GTP is catalyzed by a GTPase activating protein (GAP). The Ragulator complex functions as a GEF and promotes the active GTP-bound form. The GATOR1 complex functions as a GAP and stimulates RRAGA GTPase activity to turn it into its inactive GDP-bound form, preventing mTORC1 recruitment and activation.
Pathway
  • Instruction: What is the role of ¡protein¿ in the metabolic pathway? (Q6MUA3)
    • Response: Pyrimidine metabolism; CTP biosynthesis via de novo pathway; CTP from UDP: step 2/2.
    • Ground Truth: Pyrimidine metabolism; CTP biosynthesis via de novo pathway; CTP from UDP: step 2/2.
  • Instruction: Which metabolic pathway does ¡protein¿ associate with? (Q63AF7)
    • Response: Carbohydrate degradation; pentose phosphate pathway; D-ribose 5-phosphate from D-ribulose 5- phosphate (non-oxidative stage): step 1/1.
    • Ground Truth: Carbohydrate degradation; pentose phosphate pathway; D-ribose 5-phosphate from D-ribulose 5- phosphate (non-oxidative stage): step 1/1.
  • Instruction: What is the metabolic pathway that ¡protein¿ is involved in? (Q9SIB9)
    • Response: Carbohydrate metabolism; tricarboxylic acid cycle; isocitrate from oxaloacetate: step 2/2.; Organic acid metabolism; propanoate degradation.
    • Ground Truth: Carbohydrate metabolism; tricarboxylic acid cycle; isocitrate from oxaloacetate: step 2/2.
Subsequently, based on the above cases, we present the error patterns observed for each question type, with a comprehensive summary below:
Figure 7. Taxonomic distribution at kingdom and phylum levels in open-ended generation dataset.
  • For Function, Similarity, and Subcellular Location properties and functions, our model demonstrates high prediction accuracy due to comprehensive annotation coverage and standardized formats.
  • Regarding the three Gene Ontology-related properties and functions, the model achieves accurate predictions for the majority of proteins, with primary errors manifesting as either under-prediction or over-prediction of certain terms.
  • For other properties and functions with adequate annotation coverage (including Induction, Short Sequence Motif, Tissue Specificity, Activity Regulation, and Pathway), the model maintains prediction accuracy, with the predominant error being term omission.
  • In the case of Developmental Stage, which suffers from limited annotation coverage and poor standardization, while the model may not achieve perfect annotation accuracy, it successfully captures the essential meaning, as illustrated in this example:
    Instruction: At which specific developmental stages is <protein> expressed? (A5HEI1)
    • Response: Expressed in embryos (at protein level).
    • Ground Truth: Detected throughout the embryo, covering all stages of development from pre-globular to torpedo stages.
Table 15. Performance comparison of SEPIT-TinyLlama-MoEs across different kingdoms on OOD proteins.
Kingdom BLEU-2 ROUGE-1 ROUGE-L METEOR BERT-F1 Archaea 64.14 68.82 68.32 67.32 95.35 Bacteria 64.73 72.17 71.16 69.75 95.69 Eukaryota 52.68 61.84 59.84 59.75 94.40
Analysis Based on Taxonomic Classification
Using the open-ended generation dataset as an example, we show the distribution of kingdom and phylum in the training and test sets in Figure 7. Additionally, in Table 15, we present the performance of SEPIT-TinyLlama-MoEs across different kingdoms on the OOD proteins (considering annotation coverage differences across kingdoms, we selected the 3 most comprehensively annotated properties and functions: Function, Similarity, and Subcellular location). Intuitively, since Eukaryota have significantly more training data compared to Archaea, the model would be expected to perform better on eukaryotic proteins. However, contrary to this expectation, Archaea and Bacteria demonstrate superior performance. We attribute this phenomenon to two primary factors:
  • Data complexity differences:
    • Although Eukaryota has more training data, eukaryotic genomes and protein structures typically exhibit higher complexity.
    • Eukaryotic function descriptions tend to be more elaborate, detailed, and diverse, thereby increasing the complexity of the generation task.
  • Data quality factors: Despite the larger dataset size for Eukaryota, it encompasses more noise and heterogeneous annotations, including a higher prevalence of technical terminology and intricate biological processes, which poses greater challenges for accurate prediction.